two resistances 2 ohm and 4 ohm are connected in series and a 4 ohm is in parallel between two points find equivalent resistance.
1
Expert's answer
2020-09-21T09:41:26-0400

2 resistors in series, so equivalent resistance"=2+4=6\Omega"

The one connected in parallel, so equivalent resistance"=\frac{6\times 4}{6+4}=2.4\Omega"